What is the osmotic pressure of a solution containing 83.6 g of fructose (C6H12O6, MW=180 g/mol) in two liters of solution at 79°C?
1
Expert's answer
2020-04-23T13:06:55-0400

The osmotic pressure:



R = 8.314 J/mol K - another way to express the constant!

n = m/M = 83.6 / 180 = 0,4644 mol

T = 79 + 273.15 = 352,15 K

П = (0,4644 mol * 8.314 J/mol K * 352,15 K) / 2x10-3 m3

П = 679829,4 Pa